# The Feasibility of a Spacecraft Flyby with the Third Interstellar Object 3I/ATLAS from Earth or Mars

### v1

@@ 5 Discussion
On the other hand, Mars departures require substantially less $\Delta V$. For example, a pre-discovery flyby on 2025 March 6 (four months prior to discovery) could reach 3I/ATLAS with just $\Delta V=2.02$ km s-1 while a launch on the discovery date of 2025 July 1 would require $\Delta V\sim 4$ km s-1. These $\Delta V$ values likely fall within the maneuvering envelope of current Mars orbiters, making it possible to re-target one for a close flyby of 3I/ATLAS.
There are a variety of spacecraft currently orbiting Mars that are nearing their end of life — Mars Reconnaissance Orbiter (R. W. Zurek & S. E. Smrekar, 2007), MAVEN (B. M. Jakosky et al., 2015), Mars Odyssey (R. S. Saunders et al., 2004), Mars Express (A. Chicarro et al., 2004; A. F. Chicarro, 2007), ExoMars Trace Gas Orbiter (A. C. Vandaele et al., 2015) and others. These spacecraft collectively carry a wide range of imaging cameras, spectrometers, in situ instrumentation, and communications and operation tools. Repurposing these missions as interstellar interceptors would offer valuable scientific data, especially since they may be able to reach interstellar objects that are inaccessible from Earth. While repurposing Earth-observing missions is also possible, these spacecraft are generally de-orbited during their end-of-life phase and have less maneuvering capability, limiting their use as interstellar interceptors.
The Janus spacecraft mission (D. J. Scheeres et al., 2021), part of NASA’s SIMPLEx program, consists of two identical 12 U CubeSats, designed to fly by binary asteroids. Each carries JCam, a visible/infrared imager capable of capturing meter-scale resolution images and thermal data. Janus could have potentially performed its 10 km s-1 flyby of 3I/ATLAS if we had launched it within the first few days after discovery since it is one of the ready-to-launch spacecraft. This close encounter would allow Janus to collect detailed images and thermal data, helping us to understand 3I/ATLAS’s shape, rotation, and dust activity.
@@
We thank Shannon Curry, Adina Feinstein, Karen Meech, James Wray, Avi Loeb, Devin Hoover, and Qicheng Zhang for helpful conversations.

### v2

@@ 5 Discussion
On the other hand, Mars departures require substantially less $\Delta V$. For example, a pre-discovery flyby on 2025 March 6 (four months prior to discovery) could reach 3I/ATLAS with just $\Delta V=2.02$ km s-1 while a launch on the discovery date of 2025 July 1 would require $\Delta V\sim 4$ km s-1. These $\Delta V$ values fall within the maneuvering envelope of current Mars orbiters, making it possible to re-target one for a close flyby of 3I/ATLAS.
There are a variety of spacecraft currently orbiting Mars that are nearing their end of life — Mars Reconnaissance Orbiter (R. W. Zurek & S. E. Smrekar, 2007), MAVEN (B. M. Jakosky et al., 2015), Mars Odyssey (R. S. Saunders et al., 2004), Mars Express (A. Chicarro et al., 2004; A. F. Chicarro, 2007), ExoMars Trace Gas Orbiter (A. C. Vandaele et al., 2015) and others. These spacecraft collectively carry a wide range of imaging cameras, spectrometers, in situ instrumentation, and communications and operation tools. Repurposing these missions as interstellar interceptors would offer valuable scientific data, especially since they may be able to reach interstellar objects that are inaccessible from Earth. While repurposing Earth-observing missions is also possible, these spacecraft are generally de-orbited during their end-of-life phase and have less maneuvering capability, limiting their use as interstellar interceptors. Beyond Mars orbiters, opportunistic geometries with existing deep-space spacecraft may also enable intercepts. A. Loeb et al. (2025) propose using Juno (S. J. Bolton et al., 2017) during 3I/ATLAS’s March 2026 Jupiter approach.
The Janus spacecraft mission (D. J. Scheeres et al., 2021), part of NASA’s SIMPLEx program, consists of two identical , designed to fly by binary asteroids. Each carries JCam, a visible/infrared imager capable of capturing meter-scale resolution images and thermal data.
@@
We thank Shannon Curry, Adina Feinstein, Karen Meech, James Wray, Abraham Loeb, Adam Hibberd, Devin Hoover, and Qicheng Zhang for helpful conversations.